OpenStudy (anonymous):

(16/81) to the power -3/4 ?

OpenStudy (anonymous):

Well observe:$$16=4^2=(2^2)^2=2^4\\81=9^2=(3^2)^2=3^4$$

OpenStudy (anonymous):

the exponent has three parts denominator, that means take the fourth root numerator, that means cube it a minus sign, that means take the reciprocal

OpenStudy (ivancsc1996):

That means: \[\sqrt[4]{(\frac{ 81 }{ 16 })^{3}}\]

OpenStudy (anonymous):

ooh, thanks:D

OpenStudy (anonymous):

the fourth root of \(16\) is \(2\) because \(2^4=16\) when you cube \(\) you get \(2^3=8\) repeat the process for \(81\) then flip the result

OpenStudy (anonymous):

@ivancsc1996 fourth-root is probably a better idea to do first when computing

OpenStudy (ivancsc1996):

Yes you are right. @shanzayxox98 It's better to use:\[(\sqrt[4]{\frac{ 81 }{ 16 }})^{3}\]